OpenStudy (anonymous):

The product of the two consecutive positive integers is 1 more than the sum. Find the integers.

OpenStudy (anonymous):

the equation used to set that up would be x(x+1)=1+x+(x+1)

OpenStudy (triciaal):

let the integers be a and b then ab = a + b + 1 b = a + 1 therefore a(a + 1) = a + a+1 + 1 a^2 + a = 2a + 2 a^2 -a - 2 = 0 solve quadratic a = 2, b= 3

OpenStudy (anonymous):

yeah, I got the same thing but I wasn't sure

OpenStudy (triciaal):

good practice to always check your solution ab = a + b + 1 2*3 = 2 + 3 + 1 6 = 6 therefore 2,3 correct keep practicing